Two balls of identical mass of 1.179 kg and with identical charges placed on them hang from the ceiling on identical strings of lengths 1.110 m as shown. If the angle with respect to thee vertical that the strings form is 28.5 degrees, what is the charge on each ball? Please give your answer in micro-coulomb (i.e. do not enter a unit, but calculate your answer in micro- coulomb) Submit Answer Incorrect. Tries 4/20 Previous Tries

First Person got it wrong and they said .03 or .009
Answer #1

First find distance d

sin heta =(d/2)/l

sin28.5(d/2)/1.11m

1.11sin28.5 - d/2

{color{Red} d=1.06m}

=================

Find Tension

Consider all vertical forces

sum F_{y}=0

Tcos28.5 = mq

T = 1.179kg * 9.81/cos28·5

{color{Red} T =13.16N}

==============

Consider all Horizontal Forces

Fr = 0

F=Tsin28.5

= Tsin28.5

charges are identical

9*10^{9}*q^{2}/(1.06m)^{2}=13.16N*sin28.5

9 = 13.16 * sin 28.5 * 1.062/0 * 109)

q=2.8*10^{-5}C

ANSWER: {color{Red} q=28mu C}

====================